# Will LLMs Scaling Hit the Wall? Breaking Barriers via Distributed Resources on Massive Edge Devices

### v1

@@ A.1 AI Monopoly and Democratization
The current AI landscape is characterized by significant concentration of power among a few tech giants, primarily due to their monopoly over massive computing resources and data centers (Bommasani et al., 2021). This monopolistic trend has intensified with companies like OpenAI increasingly moving towards closed systems. While open-source alternatives like Llama (Touvron et al., 2023), Deepseek (Liu et al., 2024a) and other community-driven models have made strides towards democratization by releasing model parameters and technical reports (Zhang et al., 2024b), the gap in computational resources and data access between major AI companies and other players remains substantial and continues to widen. This disparity in resources allows tech giants to maintain their absolute dominance in determining the direction of AI development, raising concerns about AI democratization.
Edge device-based collaborative training presents a promising pathway to democratize AI development (Wu et al., 2024a). By leveraging the collective computing power of millions of edge devices, this approach could effectively challenge the existing monopolistic structure (Anderson et al., 2024b; Li et al., 2024c). This democratization of AI training through edge devices could fundamentally reshape the structure of responsibilities and authorities. If everyone can participate in training LLMs, the AI landscape could fundamentally change. Training decisions would shift from companies to communities, creating shared responsibility for model development (Brown et al., 2024). Global participation would help models reflect diverse cultural perspectives, while allowing communities to adapt models for their local needs. Furthermore, this decentralized approach could foster a more competitive and innovative AI ecosystem. When the barriers to entry for AI model training are lowered, we can expect to see a broader range of specialized models emerging (like (Li et al., 2024b; Zhang et al., 2024d; Yang et al., 2024b; Zhong et al., 2024; Anderson et al., 2024a; Wang et al., 2024c)), better suited to local needs and diverse use cases.
@@ A.2 Fairness and Incentive Mechanisms
The distributed training paradigm introduces new considerations for model fairness and bias mitigation (Yao et al., 2024; Li et al., 2020). When training occurs across diverse edge devices, the resulting models can potentially better reflect the heterogeneous nature of user populations (Wang et al., 2020). However, this approach also raises concerns about participation bias, where differences in device capabilities or user engagement could lead to underrepresentation of certain groups (Kairouz et al., 2021). To address these challenges, researchers have proposed various fairness-aware federated learning algorithms (Mohri et al., 2019) that aim to ensure equitable model performance across different demographic groups and device types (Li et al., 2020).
To sustain a distributed training ecosystem, effective incentive mechanisms are crucial for motivating user participation (Kang et al., 2019). Traditional approaches like computational resource sharing (Khan et al., 2019) and privacy-preserving reward systems (Zhan et al., 2020) have shown promise in encouraging user engagement. More innovative solutions include token-based reward systems (Zhang et al., 2024c) and reputation mechanisms (Yang et al., 2019) that compensate users for their contributions while maintaining system integrity. These incentive structures not only encourage consistent participation but also help ensure the quality of contributed training data (Feng et al., 2019), creating a sustainable ecosystem for collaborative AI development.
@@ A.3 Carbon Footprint and Energy Efficiency
The shift from centralized to distributed training offers compelling environmental benefits (Yang et al., 2024c). Traditional data centers housing large language models face significant energy challenges(Li et al., 2024d) - their high-performance GPUs require extensive cooling systems that consume 30-40% of total energy (Capozzoli & Primiceri, 2015). In contrast, FL distributes computation across edge devices like smartphones and tablets that operate at much lower temperatures and power levels, eliminating industrial cooling needs (Qiu et al., 2021).
FL also dramatically reduces data transmission energy costs. While centralized approaches require raw data transfer from millions of devices, FL only transmits lightweight model updates, substantially decreasing network energy overhead (McMahan et al., 2017a). The hardware efficiency gap is striking - edge devices like the NVIDIA Tegra X2 consume just 7.5W during training compared to 250W for data center GPUs (Smith, 2017), translating to major carbon footprint reductions, particularly for simpler models (Li et al., 2018b). By reducing reliance on power-hungry data centers and leveraging existing consumer devices, FL enables more sustainable AI development through optimized energy efficiency and minimized infrastructure needs. This combination of reduced cooling requirements, efficient hardware utilization, and optimized data handling makes FL an environmentally responsible choice for the future of AI training (Lacoste et al., 2019). As climate impact becomes increasingly critical, FL’s sustainability advantages position it as a key technology for green AI development.
@@ B.1 Data: the fuel of LLMs
In the early stages of AI, despite relatively small data scales, the importance of data was already evident. The MNIST dataset, for instance, serves as a notable example. With 60,000 training images and 10,000 test images, it provided a crucial foundation for neural network research, demonstrating the fundamental role of data in model training (LeCun et al., 1998b). As data scales expanded, the capabilities of deep learning models saw significant improvements. The emergence of ImageNet, which contains 14 million images across 21,000 synsets, revolutionized computer vision. This enabled deep learning models like AlexNet to learn complex visual features and achieve breakthrough progress in image recognition tasks, reducing error rates from 26.2% to 15.3% in the ILSVRC-2012 competition (Deng et al., 2009; Krizhevsky et al., 2012). ImageNet’s success stemmed not only from its scale but also from its high quality and diversity, laying the groundwork for subsequent large-scale data applications.
@@
With the proliferation of the internet and advances in computing power, data scales have expanded dramatically, ushering AI into an era of massive data. GPT-3, for instance, was trained on 450 billion tokens, with a carefully curated mix of data sources: Common Crawl (60%), books (16%), Wikipedia (3%), and other internet-based text (21%) (Brown et al., 2020). This massive dataset enabled GPT-3 to excel across various tasks, demonstrating the decisive role of data scale in model capabilities. Compared to early datasets like MNIST and ImageNet, GPT-3’s data scale and quality reached unprecedented heights, not only advancing natural language processing but also opening new possibilities for AI generalization.
@@
Beyond scale, data quality and diversity are crucial factors in model performance. ImageNet ensures data quality through rigorous validation, with each image verified by an average of 3.3 annotators and achieving 95% accuracy in its labels (Deng et al., 2009). This precise annotation enables models to learn accurate visual features and excel in image classification tasks. In the realm of large language models, GPT-3’s training data underwent stringent cleaning and filtering, including deduplication, quality scoring based on document length and linguistic complexity, and content filtering for inappropriate content (Brown et al., 2020). This high-quality data enables GPT-3 to generate coherent and accurate text. Furthermore, diversity is essential: ImageNet covers 1,000 object categories across various domains, while GPT-3’s training data spans multiple languages, genres, and knowledge domains, providing rich linguistic knowledge and contextual understanding.
@@
A model’s capability depends on the knowledge it extracts from data, following empirically observed scaling laws. While increasing model parameters can enhance expressive power, without sufficient data, models cannot effectively utilize these parameters. DeepMind’s research on the Chinchilla model demonstrated that under the same compute budget, a 70B parameter model trained on 1.4T tokens outperforms a 280B parameter model trained on 0.35T tokens, achieving a 30% reduction in loss while using the same compute resources (Hoffmann et al., 2022). This finding directly supports the notion that data acts as a ceiling for model capabilities. Additionally, Meta’s research shows that while Llama 2 (70B) has 70 billion parameters, its performance largely benefits from training on 2T tokens of high-quality data, with particular emphasis on academic papers, code repositories, and books that enhance its reasoning capabilities (Touvron et al., 2023). These studies emphasize data’s central role in model training and suggest that optimal model scaling requires a balanced increase in both parameters and training data.
@@ B.2 Computing power: the engine of LLMs
Since the inception of neural networks, every breakthrough in the field of AI has been driven by the continuous improvement of computational power (Thompson et al., 2020). From the early multilayer perceptron (MLP) to the widely used large language models (LLM) today, the progress in computing power has always been a key engine for advancing AI.
As the prototype of neural networks, the MLP was initially used to solve linearly separable problems (Rosenblatt, 1958). Due to its relatively low computational demand, it could run on traditional CPU environments. However, as the complexity of neural network models increased and application scenarios expanded, computational requirements gradually rose. The emergence of Convolutional Neural Networks (CNN) and Recurrent Neural Networks (RNN) marked a surge in computational demands. CNN, through convolutional operations, effectively reduced the number of parameters, enhancing the computational efficiency of image processing tasks. Classic models such as LeNet (LeCun et al., 1998a) and AlexNet (Krizhevsky et al., 2012) achieved significant results in image classification, but this also led to a surge in computational resource demands. For example, AlexNet’s victory in the 2012 ImageNet competition was made possible by using the NVIDIA GTX 580 GPU, which significantly boosted computational performance (Krizhevsky et al., 2012).
@@
With the growing scale of neural network models, GPUs gradually became indispensable computing tools (Raina et al., 2009). The parallel computing capabilities of GPUs greatly accelerated the training process of neural networks, particularly in the field of deep learning. Meanwhile, specialized hardware for deep learning, such as Tensor Processing Units (TPUs), emerged (Jouppi et al., 2017). Compared to GPUs, TPUs offer higher efficiency and lower power consumption when performing matrix operations and deep learning tasks (Wang et al., 2019), making them the preferred hardware for training large-scale neural networks.
@@
As computational resources continued to expand, the scale of neural network model training also grew. The introduction of the Transformer architecture (Vaswani et al., 2017) revolutionized the field of natural language processing (NLP), especially with the launch of models like BERT (Devlin et al., 2018) and the GPT series (Brown et al., 2020; OpenAI, 2023b), which pushed NLP technology to new heights. However, the self-attention mechanism in the Transformer architecture has a computational complexity of $O(n^{2})$, where n represents the sequence length (Vaswani et al., 2017). This means that as the model scale and sequence length increase, the required computational power grows exponentially. For example, training large language models like GPT-3 (Brown et al., 2020) and GPT-4 (OpenAI, 2023b) involves trillions of parameters and requires thousands of GPUs or TPU nodes to support the process. This immense computational demand not only places extremely high requirements on hardware, but also on computational frameworks, storage, and communication bandwidth, creating unprecedented challenges (Patterson et al., 2021).
@@
Every leap in Artificial Intelligence has been driven by computational power (Amodei & Hernandez, 2018). From multilayer perceptrons to convolutional neural networks, and the introduction of the Transformer architecture, every innovation in models has been accompanied by an explosive growth in computational needs (Thompson et al., 2020). Particularly in the era of large language models, computational power is not only the foundational tool for model training but also the core driving force behind breakthroughs in AI performance (Kaplan et al., 2020). The success of large-scale models like GPT-4 validates that AI progress almost entirely depends on the support of more powerful computational resources (Hoffmann et al., 2022).
@@ Appendix C Smartphone Data Volume Estimation
Data volume estimation per smartphone: Based on industry reports (CounterPoint, 2021), the average smartphone storage capacity reached 100 GB in 2020. To ensure a conservative estimate, we assume that only 1% of this storage capacity (equivalent to approximately 1 GB per smartphone) is actively used for data generation and storage, including local images, video information, and other types of user-generated content. This assumption aligns with baseline usage scenarios while intentionally underestimating actual data utilization.
Number of smartphones: The growth of the number of smartphone users is an important basis for estimating the total amount of data. For this, we have referred to data from market research institutions (BankMyCell, 2023), which includes trends in changes to the number of smartphone users over time.
@@
Table 1: Trends in Smartphone Shipments and Compute Power. (Data source: (Canalys, 2025)).
@@ Appendix E Small Language Model (SLM) Architectures and Training Methods
Table 2 presents a comprehensive overview of the Small Language Model (SLM) landscape, categorized by architectures and training methodologies, according to Wang et al. (2024a). The table is organized into two main categories: (I) Transformer-Based Models, which represent the dominant architecture in current SLMs, and (II) Alternative Architecture Models, which explore novel approaches to achieve efficiency. The Transformer-Based section is further divided into models pre-trained from scratch, models derived from larger LLMs through knowledge distillation, and models created through various compression techniques (pruning, quantization, etc.). The Alternative Architecture section showcases emerging approaches like State Space Models (Mamba, Hymba), recurrent architectures (RWKV, xLSTM), and traditional encoder-decoder or encoder-only designs.
@@ I.A. Pre-Trained from Scratch
PhoneLM (Yi et al., 2024) 0.5B; 1.5B Transformer Pre-training DCLM-baseline (Li et al., 2024a), StarCoderData (Li et al., 2023)
Llama 3.2 (AI, 2024) 1B; 3B Transformer Pre-training, SFT, RLHF, DPO Not released (9T tokens)
Qwen 1/1.5/2/2.5 (Yang et al., 2024a; Bai et al., 2023) 0.5B-7B Transformer Pre-training Not released
Gemma/Gemma 2 (Google Team, 2024a, b) 2B; 7B Transformer Pre-training Unknown
SmolLM (Allal et al., 2024) 135M-1.7B Transformer Pre-training SmolLM corpus (Benallal et al., 2024)
H2O-Danube3 (Pfeiffer et al., 2024) 500M; 4B Transformer Pre-training (multi-stage) Unknown
MiniCPM (Hu et al., 2024) 1.2B; 2.4B Transformer Pre-training Dolma (AI2, 2023), C4 (Raffel et al., 2020)
CT-LLM (Du & Wei, 2024) 2B Transformer Pre-training MAP-CC
OLMo (Groeneveld et al., 2024) 1B; 7B Transformer Pre-training Dolma (AI2, 2023) (multiple sources)
TinyLlama (Zhang et al., 2024a) 1B Transformer Pre-training SlimPajama (Systems & Computer, 2023)
Phi-series (Abdin et al., 2024; Javaheripi et al., 2023) 1.3B-6.6B Transformer Pre-training CodeTextBook (Gunasekar et al., 2023)
OpenELM (Mehta et al., 2024) 270M-3B Transformer Pre-training RefinedWeb (Penedo et al., 2023), PILE (Gao et al., 2020)
MobiLlama (Thawakar et al., 2024) 0.5B; 0.8B Transformer Pre-training LLM360 Amber
MobileLLM (Liu et al., 2024b) 125M; 350M Transformer Pre-training Unknown (1T tokens)

### v2

@@ A.1 AI Monopoly and Democratization
The current AI landscape is characterized by significant concentration of power among a few tech giants, primarily due to their monopoly over massive computing resources and data centers [99]. This monopolistic trend has intensified with companies like OpenAI increasingly moving towards closed systems. While open-source alternatives like Llama [10], Deepseek [11] and other community-driven models have made strides towards democratization by releasing model parameters and technical reports [100], the gap in computational resources and data access between major AI companies and other players remains substantial and continues to widen. This disparity in resources allows tech giants to maintain their absolute dominance in determining the direction of AI development, raising concerns about AI democratization.
Edge device-based collaborative training presents a promising pathway to democratize AI development [101]. By leveraging the collective computing power of millions of edge devices, this approach could effectively challenge the existing monopolistic structure [102]. This democratization of AI training through edge devices could fundamentally reshape the structure of responsibilities and authorities. If everyone can participate in training LLMs, the AI landscape could fundamentally change. Training decisions would shift from companies to communities, creating shared responsibility for model development [103]. Global participation would help models reflect diverse cultural perspectives, while allowing communities to adapt models for their local needs. Furthermore, this decentralized approach could foster a more competitive and innovative AI ecosystem. When the barriers to entry for AI model training are lowered, we can expect to see a broader range of specialized models emerging (like [104, 105, 106, 107, 108]), better suited to local needs and diverse use cases.
@@ A.2 Fairness and Incentive Mechanisms
The distributed training paradigm introduces new considerations for model fairness and bias mitigation [109, 110]. When training occurs across diverse edge devices, the resulting models can potentially better reflect the heterogeneous nature of user populations [111]. However, this approach also raises concerns about participation bias, where differences in device capabilities or user engagement could lead to underrepresentation of certain groups [93]. To address these challenges, researchers have proposed various fairness-aware federated learning algorithms [112] that aim to ensure equitable model performance across different demographic groups and device types [110].
To sustain a distributed training ecosystem, effective incentive mechanisms are crucial for motivating user participation [113]. Traditional approaches like computational resource sharing [114] and privacy-preserving reward systems have shown promise in encouraging user engagement. More innovative solutions include token-based reward systems [115] and reputation mechanisms [116] that compensate users for their contributions while maintaining system integrity. These incentive structures not only encourage consistent participation but also help ensure the quality of contributed training data [117], creating a sustainable ecosystem for collaborative AI development.
@@ A.3 Carbon Footprint and Energy Efficiency
The shift from centralized to distributed training offers compelling environmental benefits [118]. Traditional data centers housing large language models face significant energy challenges[119] - their high-performance GPUs require extensive cooling systems that consume 30-40% of total energy [120]. In contrast, FL distributes computation across edge devices like smartphones and tablets that operate at much lower temperatures and power levels, eliminating industrial cooling needs [121].
FL also dramatically reduces data transmission energy costs. While centralized approaches require raw data transfer from millions of devices, FL only transmits lightweight model updates, substantially decreasing network energy overhead [122]. The hardware efficiency gap is striking - edge devices like the NVIDIA Tegra X2 consume just 7.5W during training compared to 250W for data center GPUs [123], translating to major carbon footprint reductions, particularly for simpler models [124]. By reducing reliance on power-hungry data centers and leveraging existing consumer devices, FL enables more sustainable AI development through optimized energy efficiency and minimized infrastructure needs. This combination of reduced cooling requirements, efficient hardware utilization, and optimized data handling makes FL an environmentally responsible choice for the future of AI training [125]. As climate impact becomes increasingly critical, FL’s sustainability advantages position it as a key technology for green AI development.
@@ B.1 Data: the fuel of LLMs
In the early stages of AI, despite relatively small data scales, the importance of data was already evident. The MNIST dataset, for instance, serves as a notable example. With 60,000 training images and 10,000 test images, it provided a crucial foundation for neural network research, demonstrating the fundamental role of data in model training [8]. As data scales expanded, the capabilities of deep learning models saw significant improvements. The emergence of ImageNet, which contains 14 million images across 21,000 synsets, revolutionized computer vision. This enabled deep learning models like AlexNet to learn complex visual features and achieve breakthrough progress in image recognition tasks, reducing error rates from 26.2% to 15.3% in the ILSVRC-2012 competition [9, 126]. ImageNet’s success stemmed not only from its scale but also from its high quality and diversity, laying the groundwork for subsequent large-scale data applications.
@@
With the proliferation of the internet and advances in computing power, data scales have expanded dramatically, ushering AI into an era of massive data. GPT-3, for instance, was trained on 450 billion tokens, with a carefully curated mix of data sources: Common Crawl (60%), books (16%), Wikipedia (3%), and other internet-based text (21%) [5]. This massive dataset enabled GPT-3 to excel across various tasks, demonstrating the decisive role of data scale in model capabilities. Compared to early datasets like MNIST and ImageNet, GPT-3’s data scale and quality reached unprecedented heights, not only advancing natural language processing but also opening new possibilities for AI generalization.
@@
Beyond scale, data quality and diversity are crucial factors in model performance. ImageNet ensures data quality through rigorous validation, with each image verified by an average of 3.3 annotators and achieving 95% accuracy in its labels [9]. This precise annotation enables models to learn accurate visual features and excel in image classification tasks. In the realm of large language models, GPT-3’s training data underwent stringent cleaning and filtering, including deduplication, quality scoring based on document length and linguistic complexity, and content filtering for inappropriate content [5]. This high-quality data enables GPT-3 to generate coherent and accurate text. Furthermore, diversity is essential: ImageNet covers 1,000 object categories across various domains, while GPT-3’s training data spans multiple languages, genres, and knowledge domains, providing rich linguistic knowledge and contextual understanding.
@@
A model’s capability depends on the knowledge it extracts from data, following empirically observed scaling laws. While increasing model parameters can enhance expressive power, without sufficient data, models cannot effectively utilize these parameters. DeepMind’s research on the Chinchilla model demonstrated that under the same compute budget, a 70B parameter model trained on 1.4T tokens outperforms a 280B parameter model trained on 0.35T tokens, achieving a 30% reduction in loss while using the same compute resources [2]. This finding directly supports the notion that data acts as a ceiling for model capabilities. Additionally, Meta’s research shows that while Llama 2 (70B) has 70 billion parameters, its performance largely benefits from training on 2T tokens of high-quality data, with particular emphasis on academic papers, code repositories, and books that enhance its reasoning capabilities [10]. These studies emphasize data’s central role in model training and suggest that optimal model scaling requires a balanced increase in both parameters and training data.
@@ B.2 Computing power: the engine of LLMs
Since the inception of neural networks, every breakthrough in the field of AI has been driven by the continuous improvement of computational power [127]. From the early multilayer perceptron (MLP) to the widely used large language models (LLM) today, the progress in computing power has always been a key engine for advancing AI.
As the prototype of neural networks, the MLP was initially used to solve linearly separable problems [128]. Due to its relatively low computational demand, it could run on traditional CPU environments. However, as the complexity of neural network models increased and application scenarios expanded, computational requirements gradually rose. The emergence of Convolutional Neural Networks (CNN) and Recurrent Neural Networks (RNN) marked a surge in computational demands. CNN, through convolutional operations, effectively reduced the number of parameters, enhancing the computational efficiency of image processing tasks. Classic models such as LeNet [129] and AlexNet [126] achieved significant results in image classification, but this also led to a surge in computational resource demands. For example, AlexNet’s victory in the 2012 ImageNet competition was made possible by using the NVIDIA GTX 580 GPU, which significantly boosted computational performance [126].
@@
With the growing scale of neural network models, GPUs gradually became indispensable computing tools [130]. The parallel computing capabilities of GPUs greatly accelerated the training process of neural networks, particularly in the field of deep learning. Meanwhile, specialized hardware for deep learning, such as Tensor Processing Units (TPUs), emerged [131]. Compared to GPUs, TPUs offer higher efficiency and lower power consumption when performing matrix operations and deep learning tasks [132], making them the preferred hardware for training large-scale neural networks.
@@
As computational resources continued to expand, the scale of neural network model training also grew. The introduction of the Transformer architecture [39] revolutionized the field of natural language processing (NLP), especially with the launch of models like BERT [3] and the GPT series [5, 4], which pushed NLP technology to new heights. However, the self-attention mechanism in the Transformer architecture has a computational complexity of $O(n^{2})$, where n represents the sequence length [39]. This means that as the model scale and sequence length increase, the required computational power grows exponentially. For example, training large language models like GPT-3 [5] and GPT-4 [4] involves trillions of parameters and requires thousands of GPUs or TPU nodes to support the process. This immense computational demand not only places extremely high requirements on hardware, but also on computational frameworks, storage, and communication bandwidth, creating unprecedented challenges [7].
@@
Every leap in Artificial Intelligence has been driven by computational power [133]. From multilayer perceptrons to convolutional neural networks, and the introduction of the Transformer architecture, every innovation in models has been accompanied by an explosive growth in computational needs [127]. Particularly in the era of large language models, computational power is not only the foundational tool for model training but also the core driving force behind breakthroughs in AI performance [1]. The success of large-scale models like GPT-4 validates that AI progress almost entirely depends on the support of more powerful computational resources [2].
@@ Appendix C Smartphone Data Volume Estimation
Data volume estimation per smartphone: Based on industry reports [134], the average smartphone storage capacity reached 100 GB in 2020. To ensure a conservative estimate, we assume that only 1% of this storage capacity (equivalent to approximately 1 GB per smartphone) is actively used for data generation and storage, including local images, video information, and other types of user-generated content. This assumption aligns with baseline usage scenarios while intentionally underestimating actual data utilization.
Number of smartphones: The growth of the number of smartphone users is an important basis for estimating the total amount of data. For this, we have referred to data from market research institutions [46], which includes trends in changes to the number of smartphone users over time.
@@
Table 1: Trends in Smartphone Shipments and Compute Power. (Data source: [55]).
@@ Appendix E Small Language Model (SLM) Architectures and Training Methods
Table 2 presents a comprehensive overview of the Small Language Model (SLM) landscape, categorized by architectures and training methodologies, according to [57]. The table is organized into two main categories: (I) Transformer-Based Models, which represent the dominant architecture in current SLMs, and (II) Alternative Architecture Models, which explore novel approaches to achieve efficiency. The Transformer-Based section is further divided into models pre-trained from scratch, models derived from larger LLMs through knowledge distillation, and models created through various compression techniques (pruning, quantization, etc.). The Alternative Architecture section showcases emerging approaches like State Space Models (Mamba, Hymba), recurrent architectures (RWKV, xLSTM), and traditional encoder-decoder or encoder-only designs.
@@ I.A. Pre-Trained from Scratch
PhoneLM [135] 0.5B; 1.5B Transformer Pre-training DCLM-baseline [136], StarCoderData [137]
Llama 3.2 [138] 1B; 3B Transformer Pre-training, SFT, RLHF, DPO Not released (9T tokens)
Qwen 1/1.5/2/2.5 [139, 140] 0.5B-7B Transformer Pre-training Not released
Gemma [141] 2B; 7B Transformer Pre-training Unknown
SmolLM2 [142] 135M-1.7B Transformer Pre-training SmolLM corpus [143]
H2O-Danube3 [144] 500M; 4B Transformer Pre-training (multi-stage) Unknown
MiniCPM [145] 1.2B; 2.4B Transformer Pre-training Dolma [146], C4 [14]
CT-LLM [147] 2B Transformer Pre-training MAP-CC
OLMo [148] 1B; 7B Transformer Pre-training Dolma [146] (multiple sources)
TinyLlama [149] 1B Transformer Pre-training SlimPajama [65]
Phi-series [150, 151] 1.3B-6.6B Transformer Pre-training CodeTextBook [152]
OpenELM [153] 270M-3B Transformer Pre-training RefinedWeb [154], PILE [155]
MobiLlama [156] 0.5B; 0.8B Transformer Pre-training LLM360 Amber
MobileLLM [157] 125M; 350M Transformer Pre-training Unknown (1T tokens)